What is the solution to the following system of equations?
2x+3y=6
x+y=4

Answer by checkley77(12844)

2x+3y=6
x+y=4 MULTIPLY THIS EQUATION BY -2 & ADD.
-2X-2Y=-8
2X+3Y=6
--------------
Y=-2 ANS.
X-2=4
X=4+2
X=6 ANS.
PROOF:
2*6+3*-2=6
12-6=6
6=6